OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (anonymous):

I don't want anybody to do it for me just am stuck. could use a little help!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (tkhunny):

It is good that you are stuck, since there isn't a question. All we have is a description of the situation.

OpenStudy (anonymous):

whats the question ?

OpenStudy (anonymous):

1. Change the equation into slope-intercept form . Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

2x + 3y = 1200 slope intercept form is y = mx + b 3y = -2x + 1200 y = -2/3x + 1200/3 y = -2/3x + 400 ==> slope intercept form In y = mx + b form, the slope is in the m position and the y intercept is in the b position. Therefore, the slope is -2/3 and the y intercept is 400

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

thank you and no

OpenStudy (anonymous):

hey hold on

OpenStudy (anonymous):

can you answer one more question

OpenStudy (anonymous):

possibly...what ya got

OpenStudy (anonymous):

Describe how you would graph this line using slope intercept method

OpenStudy (anonymous):

write the equation in function notation. Explain what the graph of the function reprsdents

OpenStudy (anonymous):

doesent matter which one

OpenStudy (anonymous):

@kelliegirl33 help please

OpenStudy (anonymous):

y = -2/3x + 400 we already know that the y intercept is 400 or (0,400) to find the x intercept, sub in 0 for y 2x + 3(0) = 1200 2x = 1200 x = 600...so the x intercept is 600 or (600,0) plot your intercept points (0,400) and (600,0) your slope is -2/3. So start at your x axis (600,0) and come down 2 spaces, and go to the right 3 spaces, then down 2 and to the right 3...keep doing this and you will eventually cross the y axis at (600,0)

OpenStudy (anonymous):

is that for the second or third question I asked

OpenStudy (anonymous):

that is graphing using slope intercept

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

function notation...solve for y and then replace y with f(x). y = -2/3x + 400 function notation would be : f(x) = -2/3x + 400

OpenStudy (anonymous):

ok thank you for all the help